OpenStudy (booklover40):

Let x=4 and y=-2. Evaluate(x^2-y^2(10-y^2)/ 3)^2.

OpenStudy (bobo-i-bo):

Whenever there's an x, replace it with 4, whenever there's a y, replace it with -2. Then evaluate.

OpenStudy (jdoe0001):

\(\left({\color{brown}{ x}}^2-{\color{blue}{ y}}^2\cfrac{10-{\color{blue}{ y}}^2}{3} \right)^2\qquad \begin{cases} {\color{brown}{ x}}=4\\ {\color{blue}{ y}}=-2 \end{cases}\implies \left({\color{brown}{ 4}}^2-{\color{blue}{ (-2)}}^2\cfrac{10-{\color{blue}{ (-2)}}^2}{3} \right)^2\)

rebeccaxhawaii (rebeccaxhawaii):

mhm ^^